OpenStudy (anonymous):

(5x^3-6x^2 + 5)-(5-6x^2+5x^3

Nnesha (nnesha):

do you know there is "edit" button ???

Nnesha (nnesha):

so you don't have to close and post new one :)

OpenStudy (anonymous):

didn't notice it

Nnesha (nnesha):

okay so what you got after distributing 2nd parenthesis by -1 ?

OpenStudy (solomonzelman):

A - A = 0 regardless of what "A" is (provided that it is a defined value)

OpenStudy (anonymous):

-5 -6x^2 -5x^3

Nnesha (nnesha):

what is -1 times -6x^2 = ??

OpenStudy (anonymous):

6x^2

OpenStudy (anonymous):

next step?

Nnesha (nnesha):

now combine like terms \[\huge\rm 5x^3 - 6x^2 + 5 - 5 +6x^2 -5x^3\] listen don't forget signs !!! and remember if both variables have same exponent then you can combine them otherwise nope!

OpenStudy (anonymous):

10x^6 -12x^4

Nnesha (nnesha):

wait nope just lO_Ok at there focus on signs see what magic gonna happen '

OpenStudy (anonymous):

so just 10 n 12

OpenStudy (solomonzelman):

(5x^3-6x^2 + 5)-(5-6x^2+5x^3) (5x^3-6x^2 + 5)-(5x^3-6x^2+5) just like any A-A=0 so is this =?

OpenStudy (solomonzelman):

no need distributing

Nnesha (nnesha):

nope  what is 5x^3 - 5x^3 =?